### v1

# Infinite Width Models That Work: Why Feature Learning Doesn’t Matter as Much as You Think

Abstract
Common infinite-width architectures such as Neural Tangent Kernels (NTKs) have historically shown weak performance compared to finite models. This has been attributed to the absence of feature learning. We show that this is not the case. In fact, we show that infinite width NTK models are able to access richer features than finite models by selecting relevant subfeatures from their (infinite) feature vector. In fact, we show experimentally that NTKs under-perform traditional finite models even when feature learning is artificially disabled. Instead, weak performance is due to the fact that existing constructions depend on weak optimizers like SGD. We provide an infinite width limit based on ADAM-like learning dynamics and demonstrate empirically that the resulting models erase this performance gap.
1 Introduction
Neural Tangent Kernels (NTKs) [1] represent one of the first attempts to capture the dynamics of training in the infinite width limit. Unfortunately, they often underperform finite-width models. The usual explanation given is that these infinite limits don’t allow feature learning [2].
By “features,” we mean the inputs to the last layer of a deep learning model. “Feature learning” refers to the evolution of the rest of the model in order to learn useful feature representations. This language invites us to think of the entire first part of the model as merely setting up the last layer with the right representations to solve downstream tasks linearly.
In [2], Yang and Hu prove their “Dynamical Dichotomy”, which states that under fairly general conditions, the only models that admit an infinite width kernel representation are those where the features do not evolve during training. The entire model except for the last layer is effectively frozen at initialization.
Since the bulk of practical infinite width schemes are kernel machines, this was taken as something of a death knell for efficient, expressive infinite models. How could infinite models hope to learn complex behavior if most of the model is untrained? Even a large feature vector will be mostly useless at initialization.
But “infinite” is very different from “large,” and sometimes our intuitions can lead us astray. At infinite width, feature vectors contain an infinite variety of structure, all of which is available to the final layer.
We’ll prove that merely by upweighting and downweighting subvectors, the final layer of an infinite width model has access to every possible behavior the previous layers could have learned. This can be viewed as an infinite extension to random kitchen sinks [3] approach. Our contribution is showing that at the infinite limit, the effect of this freezing vanishes altogether.
We then offer a simpler explanation for the performance gap: common infinite width limits capture the dynamics of training under stochastic gradient descent (SGD), while the most successful finite models are trained using more sophisticated optimizers like ADAM [4].
To address this issue, we provide a construction for infinite width limits under ADAM-like learning dynamics. This construction preserves the kernel representation of the limit, but eliminates the performance gap with finite models.
2 Defining an infinite model
We’ll begin by writing an ordinary deep learning model in a way that makes it easier to think about both feature learning and infinite limits. For arbitrary input space $X$ and output space $Y=\mathbb{R}^{E}$, let $\left(x_{1},y_{1}\right),\left(x_{2},y_{2}\right),\dots,\left(x_{N},y_{N}\right)\in X\times Y$ be training examples. Let $L:Y\times Y\rightarrow\mathbb{R}$ be a differentiable loss function.
Let $H$ be the number of features. This is the quantity we will ultimately push to infinity.
Let $f_{M\theta}:X\rightarrow\mathbb{R}^{E}$ be an arbitrary deep learning model, where $M\in\mathbb{R}^{H\times E}$ denotes the weight matrix of the final layer and $\theta\in\mathbb{R}^{P}$ denotes the remaining $P$ parameters. We will use $g_{\omega_{h}}:X\rightarrow\mathbb{R}$ to denote the submodel generating coordinate $h$ of the feature vector. It is parameterized by some subset of the parameters $\omega_{h}\subseteq\theta$ of size $S$. (The subsets for different $h$ need not be disjoint, but will all have the same size.)
Altogether the model is given by
$f_{M\theta}\left(x\right)=\frac{1}{\sqrt{H}}\sum^{H}_{h=1}g_{\omega_{h}}\left(x\right)M_{h}$ (1)
where we have scaled by $1/\sqrt{H}$ to keep values finite as we scale $H$.
We would like to train this model to minimize the loss over the training set,
$\min_{M,\theta}\sum^{N}_{i=1}L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right).$
Note that this is an extremely general formulation. $g$ may be any model, including models which contain their own infinite width submodels. $L$ is also arbitrary, and may include other nonlinearities, or even additional submodels of its own, which need not be frozen.
All we have really said is that our model contains a matrix $M$ whose input dimension we would like to drive to infinity.
2.1 The infinite width limit
If $\Omega$ is a discrete random variable uniformly distributed over $\left\{\omega_{1},\omega_{2},\dots,\omega_{H}\right\}\in\mathbb{R}^{S}$, we can write (1) as an expectation,
$f_{M\theta}\left(x\right)=\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)\hat{M}_{\omega}\right].$ (2)
where $\hat{M}=\sqrt{H}M$ and $\hat{M}_{\omega}\in\mathbb{R}^{E}$ denotes the row of $\hat{M}$ associated with $\omega$ (for example, $\hat{M}_{\omega_{1}}=\sqrt{H}M_{1}$).
Written in this form, we could just as easily take $\Omega$ to be a continuous random variable. (In this view, $\hat{M}$ is a possibly-nonlinear map from any $\omega$ in the support of $\Omega$ to a corresponding vector in $\mathbb{R}^{E}$).
This notation has the convenient property of unifying over the finite and infinite cases depending on the choice of $\Omega$. We may define $\Omega$ over a set of any size, finite or infinite. (We haven’t yet shown that this expectation (and $\hat{M}$ in particular) is finite, but once we have written it as a kernel machine, this will be apparent.)
2.2 Infinite width models are the expected value of finite ones
Intriguingly, (2) has another, equally important meaning. Let $\omega_{h}$ be initialized from $\Omega$. Then from (1), the expected value of the model (taken over the random initialization) will be
$\mathop{\mathbb{E}}\left[\frac{1}{\sqrt{H}}\sum^{H}_{h=1}g_{\omega_{h}}\left(x\right)M_{h}\right]=\frac{1}{H}\sum^{H}_{h=1}\mathop{\mathbb{E}}_{\omega_{h}\sim\Omega}\left[g_{\omega_{h}}\left(x\right)\hat{M}_{h}\right]=\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)\hat{M}_{\omega}\right]$
This is exactly the expression in (2). In other words, the expected value of a finite model is an infinite one.
This observation makes it possible to apply results about infinite models to finite ones. In the development that follows, we encourage the reader to consider the results from both of these perspectives.
3 The Neural Tangent Kernel
We’d like to express our infinite limit in a form that makes it tractable to compute. Neural Tangent Kernels (NTKs) do this by rewriting it as a kernel machine. We review the development of this kernel. The construction given here is a simpler variant of the original [1].
Dynamical Dichotomy shows that NTK representations are not possible without essentially freezing the first part of the model. In the original development of NTKs [1], this freezing happens as an indirect (and presumably unintended) consequence of the parameterization. Here we take the simpler path of taking $\theta$ to be frozen explicitly.
We’ll make a few other simplifying assumptions, namely that we are training with batch size $1$ and that $M$ is zero initialized. These are purely cosmetic assumptions to reduce bookkeeping.
3.1 A simple NTK
After $t$ steps of training under SGD with step size $\alpha$, the value of $M$ is just the sum of the its updates under SGD.
$M_{h}$ $=-\alpha\sum^{t}_{i=1}\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial M_{h}}$
$=-\alpha\frac{1}{\sqrt{H}}\sum^{t}_{i=1}g_{\omega_{h}}\left(x_{i}\right)\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}$
where the second step follows by direct computation of the matrix derivative. It follows that $\hat{M}_{\omega}$ is finite (as we hoped) and given by
$\hat{M}_{\omega}=-\alpha\sum^{t}_{i=1}g_{\omega}\left(x_{i}\right)\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}.$ (3)
Plugging this into (2), we get
$f_{M\theta}\left(x\right)=-\alpha\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)\sum^{t}_{i=1}g_{\omega}\left(x_{i}\right)\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}\right].$
By changing the order of the summation, this can be written as
$f_{M\theta}\left(x\right)=-\alpha\sum^{t}_{i=1}\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)g_{\omega}\left(x_{i}\right)\right]\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}.$
The expectation over $g_{\omega}\left(x\right)g_{\omega}\left(x_{i}\right)$ is known as the Neural Tangent Kernel (NTK) of the model, often written as $\Theta\left(x,x_{i}\right)=\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)g_{\omega}\left(x_{i}\right)\right]$. Since $\theta$ (and therefore $\Omega$) is frozen, $\Theta$ is stationary throughout training.
Again, notice that because $\Omega$ is permitted to be a discrete distribution, the NTK is defined for both finite and infinite models.
In the infinite-width limit, many common choices of $g$ give rise to a $\Theta$ that has a closed-form representation [5][6], making it tractable to compute exactly. Altogether, our kernel machine is
$f_{M\theta}\left(x\right)=-\alpha\sum^{t}_{i=1}\Theta\left(x,x_{i}\right)\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}.$ (4)
Note that unlike the standard development [1] of NTKs, this kernel is compact relative to the size of the model, having only $H$ terms in the summation, as opposed to $HE+P$ in the original. It is also stationary even in the finite case. These properties are possible because we took $\theta$ to be frozen. Dynamical Dichotomy shows that this is implicitly true already for infinite-width NTKs, all we have done is made it explicit.
4 Feature Learning for Neural Tangent Kernels
Everything we’ve done so far assumes that $\theta$ is frozen. But this kind of aggressive freezing intuitively seems like it would damage the model’s ability to learn.
Perhaps surprisingly, we show that in the infinite-width limit, freezing the model this way has no effect at all on the expressiveness of the model. The reason for this is related to the lottery ticket hypothesis [7].
In large models, it is often the case that a subset of the parameters “get lucky” and end up close to a useful value. During training, the model can improve by simply pruning the unlucky ones. This effect becomes more pronounced as the model size increases, until in the continuous limit every useful feature representation (and every useless one) exists with probability one. The final layer is free to simply select the subsets of the feature vector that are worthwhile, “training” the features without touching their weights. Unlike in the finite case, this filtering process need not reduce the cardinality of the features. There are still infinitely many.
We show that this strategy is able to entirely replace gradient descent based feature learning.
Theorem 1.
For any infinite model as in (2), suppose $\Omega$ has a differentiable density. Then we may replace arbitrary updates to $\Omega$ with appropriate updates to $\hat{M}$ without changing model behavior or learning dynamics. It follows that we can freeze $\Omega$ without affecting the final model quality.
Proof.
Since $\Omega$ has a differentiable density $\rho$, we may write (2) as
[…]
Suppose we want to update $\Omega$ by having its density $\rho$ flow along some vector field $V\left(\omega\right)$, scaled by a learning rate $\alpha$. (In SGD, we would take $V$ to be proportional to the negative gradient, $V\left(\omega\right)=-\frac{\partial L}{\partial\omega}$, but our discussion will consider the general case.) How does this update affect our model?
Note that since $V$ is arbitrary, we may break large updates into several small steps. This means that without loss of generality we make take $\alpha$ to be small.
The continuity equation [8] then tells us that if the change in $\Omega$ is $V$ the change in the density $\rho$ will be
$\Delta\rho$ $=-\alpha\nabla\cdot\left(\rho V\right)$
$=-\alpha\left(\rho\nabla\cdot V+V\cdot\nabla\rho\right)$
$=-\alpha\rho\left(\nabla\cdot V+V\cdot\nabla\log\rho\right)$
via the product rule and the score function trick.
After the update, the new model output $f^{\prime}_{M\theta}$ can be computed as
[…]
where $\hat{M}^{\prime}_{\omega}=\hat{M}_{\omega}\bigg{(}1-\alpha\big{(}\nabla\cdot V+V\cdot\nabla\log\rho\left(\omega\right)\big{)}\bigg{)}$. In other words, we can update $\hat{M}$ rather than $\rho$. ∎
This gives a procedure for converting updates to the whole model into equivalent updates on $\hat{M}$ alone.
One might legitimately object that updating $\hat{M}$ in this fashion is different from the update we get from running SGD on $\hat{M}$, and therefore can’t necessarily be found via gradient descent. But this change is a small perturbation of $\hat{M}$. If it is beneficial (i.e., having lower loss), it will be in the direction of descent.
5 Infinite width learning with ADAM
If feature learning isn’t the problem, why aren’t NTK models more competitive with conventional models? The answer is simpler than it might seem. NTK optimization has historically been designed to mimic the update performed in fixed learning rate SGD. The derivation assumes that gradients are being applied directly. But modern machine learning is almost exclusively done using optimizers that support, at a minimum, momentum. We give an NTK construction that models ADAM-style gradient updates. As we’ll see, this closes the performance gap.
Definition 1 (ADAM).
Recall that for a given gradient $J_{t}$, the ADAM update $u_{t}$ for training step $t$ can be calculated as
$m_{0}$ $=0$
$v_{0}$ $=0$
$m_{t}$ $=\beta_{1}m_{t-1}+\left(1-\beta_{1}\right)J_{t}$
$v_{t}$ $=\beta_{2}v_{t-1}+\left(1-\beta_{2}\right)J_{t}^{2}$
$u_{t}$ $=-\alpha\frac{m_{t}}{1-\beta_{1}^{t}}\cdot\left(\frac{v_{t}}{1-\beta_{2}^{t}}\right)^{-1/2}$
where $\beta_{1}$ and $\beta_{2}$ are decay hyperparameters, and all operations are applied coordinatewise.
Unfortunately, we can’t construct a NTK infinite limit based on this definition as-is. $J_{t}$, $m_{t}$, and $v_{t}$ are all infinite. In the case of SGD, we were able to find a rewriting that meant we never had to write down $J_{t}$ explicitly.
Getting a similar rewriting to work for ADAM requires a minor modification. We introduce ADAM*, a variant that is built to preserve the useful learning dynamics of ADAM while accommodating infinite limits.
Definition 2 (ADAM*).
For a given gradient $J_{t}$, the ADAM* update $u_{t}$ is the same as ADAM, except that $v_{t}$ is replaced with its expectation over $\Omega$. (For a finite model, this corresponds to averaging over any axis of dimension $H$.)
$m_{0}$ $=0$
$v_{0}$ $=0$
$m_{t}$ $=\beta_{1}m_{t-1}+\left(1-\beta_{1}\right)J_{t}$
$v_{t}$ $=\beta_{2}v_{t-1}+\left(1-\beta_{2}\right)\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[J_{t}^{2}\right]$
$u_{t}$ $=-\alpha\frac{m_{t}}{1-\beta_{1}^{t}}\cdot\left(\frac{v_{t}}{1-\beta_{2}^{t}}\right)^{-1/2}$
where the coordinatewise multiplication in the last step is taken to be broadcast (as it is now between a $H\times E$ matrix and a $E$-dimensional vector).
This change is enough to allow us to build an NTK that follows ADAM* dynamics.
Theorem 2.
Let $f_{M\theta}$ be defined as in (2). Let $M$ be zero-initialized and trained via ADAM* for $t$ steps. Let $\Theta\left(x,x_{i}\right)=\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)g_{\omega}\left(x_{i}\right)\right]$ be the NTK. Then
$f_{M\theta}\left(x\right)=-\alpha\sum^{t}_{i=1}\Theta\left(x,x_{i}\right)\frac{\partial L\left(f_{M\theta}\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}c_{i}.$ (5)
where $c_{i}=\sum^{t}_{j=i}\beta_{1}^{j-i}\frac{1-\beta_{1}}{1-\beta_{1}^{j}}\left(\frac{\hat{v}_{j}}{1-\beta_{2}^{j}}\right)^{-1/2}$ and $\hat{v}_{j}$ is defined recursively as
$\hat{v}_{0}=0,\hskip 30.00005pt\hat{v}_{j}=\beta_{2}\hat{v}_{j-1}+\left(1-\beta_{2}\right)\Theta(x_{j},x_{j})\left(\frac{\partial L\left(f_{M\theta}\left(x_{j}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{j}\right)}\right)^{2},$
with all operations performed coordinatewise.
Note that $c_{i}$ is an $E$-dimensional vector whose scale is independent of $H$. Provided we have an $O(1)$ way to compute $\Theta$, this entire computation may be performed in $O(t)$ time, even in the infinite limit. Just as for traditional NTKs, the computation is defined regardless of the scale, finite or infinite.
But unlike traditional NTKs, this NTK captures the momentum and adaptive step size that make ADAM so effective.
6 Experimental Results
Here we show experimentally that NTK models exhibit a performance gap, even when compared against models with no feature learning. We show that using the ADAM* NTK from Definition 2 closes this gap.
Our model is a six layer, decoder-only transformer model [9] with an embed dimension of 512, trained on C4 for 40k steps. We replace the MLPs with one of a few possibilities.
  • An MLP of the same size with no feature learning (i.e. frozen first matrix).
  • A (finite) NTK of the same width, using the standard construction.
  • An NTK of the same width using the ADAM* construction (ours).
  • The original MLP, unfrozen.
Figure 1: Final loss at various MLP widths for different models and NTKs.
Notice that the traditional NTK underperforms an ordinary MLP even when the MLP has no feature learning. This demonstrates that feature learning cannot be the root cause of the performance drop.
On the other hand, our ADAM* NTK has access to momentum and adaptive learning rate, and is able to closely track the performance of the frozen feature model across a wide variety of scales. In the infinite limit, our theoretical analysis suggests that this should ultimately converge with the red line, but for performance reasons we leave this to future work.
Appendix A Appendix
Proof of Theorem 2.
We’ll first prove that $\hat{v}$ as defined in Theorem 2 is related to $v$ in Definition 2. Specifically, we’ll show that $\hat{v}=Hv$.
Since $J$ is the gradient of the loss with respect to $M$, the value of row $h$ at time $t$ will be
$J_{t\omega_{h}}=\frac{\partial L\left(f_{M}\theta\left(x_{t}\right),y_{t}\right)}{\partial M_{h}}=\frac{1}{\sqrt{H}}g_{\omega_{h}}\left(x_{t}\right)\frac{\partial L\left(f_{M}\theta\left(x_{t}\right),y_{t}\right)}{\partial f_{M\theta}\left(x_{t}\right)}.$
It follows that
$\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[J_{t\omega}^{2}\right]$ $=\frac{1}{H}\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x_{t}\right)^{2}\left(\frac{\partial L\left(f_{M}\theta\left(x_{t}\right),y_{t}\right)}{\partial f_{M\theta}\left(x_{t}\right)}\right)^{2}\right]$
$=\frac{1}{H}\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x_{t}\right)^{2}\right]\left(\frac{\partial L\left(f_{M}\theta\left(x_{t}\right),y_{t}\right)}{\partial f_{M\theta}\left(x_{t}\right)}\right)^{2}$
$=\frac{1}{H}\Theta\left(x_{t},x_{t}\right)\left(\frac{\partial L\left(f_{M}\theta\left(x_{t}\right),y_{t}\right)}{\partial f_{M\theta}\left(x_{t}\right)}\right)^{2}.$
The result then follows by induction on the definition.
We turn our attention to the momentum $m$. Adding up all terms from 00 to $j$,
$m_{j}$ $=\left(1-\beta_{1}\right)\sum_{i=1}^{j}\beta_{1}^{j-i}J_{i}.$
$M$ is zero-initialized, so it’s equal to the sum of its updates, $M=\sum_{j=1}^{t}u_{j}$. $\hat{M}$ is then
$\hat{M}$ $=\sqrt{H}M$
$=\sqrt{H}\sum_{j=1}^{t}u_{j}$
$=-\alpha\sqrt{H}\sum_{j=1}^{t}\frac{m_{j}}{1-\beta_{1}^{j}}\cdot\left(\frac{v_{j}}{1-\beta_{2}^{j}}\right)^{-1/2}$
$=-\alpha\sum_{j=1}^{t}\frac{m_{j}}{1-\beta_{1}^{j}}\cdot\left(\frac{\hat{v}_{j}}{1-\beta_{2}^{j}}\right)^{-1/2}$
$=-\alpha\sum_{j=1}^{t}\left(\left(1-\beta_{1}\right)\sum_{i=1}^{j}\beta_{1}^{j-i}J_{i}\right)\frac{1}{1-\beta_{1}^{j}}\cdot\left(\frac{\hat{v}_{j}}{1-\beta_{2}^{j}}\right)^{-1/2}$
$=-\alpha\sum_{i=1}^{t}J_{i}\sum_{j=i}^{t}\beta_{1}^{j-i}\frac{1-\beta_{1}}{1-\beta_{1}^{j}}\cdot\left(\frac{\hat{v}_{j}}{1-\beta_{2}^{j}}\right)^{-1/2}$
$=-\alpha\sum_{i=1}^{t}J_{i}c_{i}$
where in the second to last step we exchange the order of summation.
From here, the proof proceeds much as it did for the ordinary NTK. From (2) we have
$f_{M\theta}\left(x\right)$ $=\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)\hat{M}_{\omega}\right]$
$=-\alpha\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)\sum_{i=1}^{t}J_{i\omega}c_{i}\right]$
$=-\alpha\sum_{i=1}^{t}\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)J_{i\omega}c_{i}\right]$
$=-\alpha\sum_{i=1}^{t}\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)g_{\omega}\left(x_{i}\right)\frac{\partial L\left(f_{M}\theta\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}c_{i}\right]$
$=-\alpha\sum_{i=1}^{t}\mathop{\mathbb{E}}_{\omega\sim\Omega}\left[g_{\omega}\left(x\right)g_{\omega}\left(x_{i}\right)\right]\frac{\partial L\left(f_{M}\theta\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}c_{i}$
$=-\alpha\sum_{i=1}^{t}\Theta\left(x,x_{i}\right)\frac{\partial L\left(f_{M}\theta\left(x_{i}\right),y_{i}\right)}{\partial f_{M\theta}\left(x_{i}\right)}c_{i},$
as desired. Note that this is essentially the same development as before, except for the introduction of $c_{i}$.